### v1

@@ Abstract
Functional brain connectome is crucial for deciphering the neural mechanisms underlying cognitive functions and neurological disorders. Graph deep learning models have recently gained tremendous popularity in this field. However, their actual effectiveness in modeling the brain connectome remains unclear. In this study, we re-examine graph deep learning models based on four large-scale neuroimaging studies encompassing diverse cognitive and clinical outcomes. Surprisingly, we find that the message aggregation mechanism, a hallmark of graph deep learning models, does not help with predictive performance as typically assumed, but rather consistently degrades it. To address this issue, we propose a hybrid model combining a linear model with a graph attention network through dual pathways, achieving robust predictions and enhanced interpretability by revealing both localized and global neural connectivity patterns. Our findings urge caution in adopting complex deep learning models for functional brain connectome analysis, emphasizing the need for rigorous experimental designs to establish tangible performance gains and perhaps more importantly, to pursue improvements in model interpretability.
@@ 1 Introduction
Brain connectome analysis has captivated researchers for its potential to decode the intricate organizations of the human brain and predict clinical outcomes [1, 2, 3, 4]. Functional magnetic resonance imaging (fMRI) plays a crucial role in this field by capturing brain activity through blood-oxygen-level-dependent (BOLD) signals. These signals facilitate the construction of functional brain networks, where nodes correspond to specific brain regions of interest (ROIs) and edges represent pair-wise statistical relationships, typically correlations, between the BOLD signals of these regions [5, 6, 7, 8]. Studying the functional connectome can enhance our understanding of brain functions [9, 10], enabling us to identify neural circuitry patterns associated with neurological conditions [11, 12] and develop improved diagnostic and therapeutic strategies [13, 14, 15, 16].
@@
Machine learning (ML) models have long been employed in functional brain connectome analysis, showing strong robustness and interpretability in linking brain connectivity to cognitive functions and neurological disorders [17, 12, 18, 19, 20, 21, 22]. These models use the correlation strength between ROIs as predictive features, without considering the underlying graph structure. Recently, graph deep learning (GDL) models like graph neural networks (GNNs) [3, 23, 24, 25, 26] and graph transformers (GTs) [4, 27, 28] have gained popularity in this field. They explicitly model brain networks as graphs, aggregating messages among nodes (ROIs) along edges (functional connections) [29]. Among various node features, the connection profile—rows of the functional connection matrix—is the most common approach [23, 25, 4, 24, 3, 26, 27, 28, 30]. Despite significant efforts to develop GDL models for brain connectome analysis, there is unfortunately no rigorous study and clear understanding of the genuine effectiveness of these advanced models over classical ML models, particularly regarding their unique message aggregation mechanisms.
@@
In this work, we conduct a comprehensive evaluation of GDL against classical ML models, across four fMRI-based brain connectome datasets: ABIDE [31], PNC [32], HCP [33], and ABCD [34]. These datasets cover a diverse population and clinical prediction tasks, from demographic attributes to cognitive abilities and neurological disorders. The experimental results are surprising: (1) classical ML models, such as logistic regression and kernel ridge regression, along with simple multi-layer perceptrons (MLPs), often match or even exceed the predictive performance of more complex GDL models; and (2) the message aggregation mechanisms unique to GDL models and believed to be a key technical advantage, appear to have a consistent detrimental effect on prediction performance within the context of functional brain networks. These results challenge prevailing assumptions about the superiority of GDL methods in modeling brain connectome.
@@
Motivated by the effectiveness of classical models and the limitations of GDL, we introduce a simple yet effective model featuring a novel dual-pathway design, which synergizes linear modeling (LM) with a graph attention network (GAT) [35]. The LM pathway efficiently captures global connectivity patterns from whole connectivity matrices, while the GAT pathway adaptively extracts local graph structures using embeddings of BOLD time series as node features. Despite its simple architecture, experimental results demonstrate that our model consistently achieves robust and competitive performance across diverse datasets, rivaling or surpassing the best-performing baselines. Beyond predictive performance, the model offers more holistic and complementary interpretability through its dual-pathway design: the GAT pathway is particularly effective at highlighting localized subnetworks and functional hubs, offering insights into modularized brain organization, while the LM pathway emphasizes global connectivity patterns and network-wide efficiency. The patterns identified not only align well with established neuroscientific knowledge but also reveal specific functional connections, salient ROIs, and subnetworks associated with cognitive processes, potentially providing finer-grained insights into neural mechanisms.
@@
The findings of this study call for careful reflection on functional brain connectome analysis, especially in using complex GDL models. They underscore the importance of thoroughly benchmarking new models in future brain connectome research. Furthermore, we advocate for future research to prioritize interpretability and contextual knowledge for scientific or clinical applications over simply pursuing prediction accuracy.
(a) Peformance on ABIDE
(b) Peformance on PNC
(c) Peformance on HCP
(d) Peformance on ABCD
@@
Figure 1: Performance comparison across different baselines on four datasets. For the ABIDE and PNC datasets, the prediction tasks are binary classification for autism disease and biological gender, respectively, evaluated using AUROC scores. For the HCP and ABCD datasets, the prediction task is fluid intelligence score prediction, a regression task evaluated using the Pearson correlation between predicted and actual values. The baselines are grouped into three categories, distinguished by color: green for graph deep learning, blue for non-graph deep learning, and red for classical machine learning models. Models highlighted with bold borders represent those with the best predictive performance among all the baselines. There is no significant difference ($p>0.05$) in performance among these highlighted models, showing that simpler models such as Logistic Regression, MLP, ElasticNet, and Kernel Ridge Regression perform comparably to the most advanced graph deep learning models. The significance bars marked with asterisks (*) indicate that the performance of non-graph deep learning models is significantly better than that of the graph deep learning models displayed to the left ($p<0.05$). Hollow circles and numbers in each box indicate the average results from 10 independent runs.
@@ 2 Results
To address the ambiguity regarding the effectiveness of complex GDL models, we conduct extensive experiments across four fMRI-based brain network datasets with various clinical prediction tasks, ranging from demographics to cognitive ability and neural disorders: ABIDE [31] for autism disease classification, PNC [32] for gender prediction, and both HCP [33] and ABCD [34] for fluid intelligence prediction. Our benchmark spans three types of models: GDL, non-graph deep learning, and classical ML models.
Fig. 1 illustrates the prediction performance of each baseline across various datasets. Simpler models often match or outperform more advanced GDL models. Specifically, for classification tasks on ABIDE and PNC datasets, classical models (e.g., Logistic Regression) and MLP-based models consistently outperform most GDL models, with performance comparable to state-of-the-art models like NeuroGraph and BrainNetTF. Similarly, for regression tasks on HCP and ABCD datasets, classical ML models such as ElasticNet [21] and Kernel Ridge Regression [12] demonstrate superior performance over most GDL models, and perform comparably to BrainNetTF [4], one of the most advanced aggregation-based GDL models. These findings challenge the prevailing belief that graph-based modeling of brain connectome is superior especially regarding the prediction performance of clinical outcomes, leading us to rethink the effectiveness of GDL models for brain connectome analysis.
(a) Varying Graph Density on ABIDE
(b) Varying Graph Density on PNC
(c) Varying Graph Density on HCP
(d) Varying Graph Density on ABCD
@@
In GDL models, the message aggregation is crucial for combining node features from neighboring nodes to form new node representations. This process is widely adopted in many fields, such as social network analysis [36], chemistry for molecular graph predictions [37], and recommendation systems [38], due to its ability to capture complex relational information. However, as shown in Fig. 1, most aggregation-based GDL models tend to underperform classical ML models like Logistic Regression, SVM/SVR [39] and Kernel Ridge Regression, and simpler MLP-based models. This raises the question about the genuine effectiveness of the aggregation operations in brain connectome analysis.
@@
To further investigate the impact of aggregation, we vary the graph density by retaining different percentages of the top $K\%$ edges in the graphs, so as to explicitly control the extent of aggregation. Notably, with $K=0$, the graphs have no edges between ROIs, so node features are transformed independently, eliminating any aggregation effects. To align with existing studies, the original settings of each GDL model are followed for extracting the top $K\%$ edges. Specifically, for models like GCN [40], GAT [35], GIN [41], GraphSage [42], NeuroGraph [26], and BrainGNN [3], only positive correlations between ROIs are considered as potential edges. This approach adheres to the common assumption that positive correlations are typically more informative in the context of functional brain networks [3, 26]. Conversely, for BrainGB [23] and BrainNetTF [4], both positive and negative correlations are incorporated, following the methodologies described in the original papers.
@@
The left panels of the Fig. 2(a)-(d) show that models like GCN, GAT, GIN, and GraphSage consistently exhibit a decline in prediction performance as the graph density increases, indicating that message aggregation does not enhance prediction accuracy and more aggregation can lead to more performance degradation in brain connectome analysis. Similarly, BrainGB and BrainGNN demonstrate optimal performance with lower density, aligning with the notion that message aggregation may not be helpful in brain connectome analysis tasks.
@@
Unlike other GDL models, BrainNetTF and NeuroGraph incorporate not only message aggregation but also residual connections, which integrate the untransformed input brain connectivity with the transformed hidden embeddings. The right panels of the Fig. 2(a)-(d) indicate that the performance of NeuroGraph and BrainNetTF remains relatively stable across varying graph density. However, when residuals are removed, as in the variants NeuroGraph w/o Res and BrainNetTF w/o Res, performance declines as density increases, consistent with other GDL models. This further underscores the point that message aggregation alone impairs predictive performance in brain connectome analysis.
@@
Previous experiments have shown that simple linear models like Logistic Regression and ElasticNet, often outperform other baselines. However, relying solely on linear models overlooks graph structure information inherent in brain networks, which is essential for understanding brain function and cognitive processes [8, 43]. To incorporate this information, it is natural to consider using graph models, which are designed to model complex relationships between nodes (ROIs) in a graph. However, Fig. 2 reveals that using connection profiles as node features diminishes the effectiveness of aggregation operations within graph models.
To address these challenges, we propose a dual-pathway model that integrates the strengths of both linear non-graph and aggregation-based graph models. The model consists of two complementary pathways. The first pathway retains the simplicity and effectiveness of linear models, which uses the flattened upper triangular part of the functional brain network matrix as input. The second pathway utilizes graph structure by embedding BOLD time series, rather than connection profiles, as node features, which are then processed by a graph attention network (GAT). Table 1 shows that this dual-pathway model outperforms baselines on the ABCD and PNC datasets, matches performance on ABIDE, and exhibits slight underperformance on HCP. Overall, this dual-pathway model achieves robust and competitive performance across four datasets.
@@
The proposed dual-pathway model integrates a non-aggregation linear model (LM) and a graph attention network (GAT) as the aggregation-based graph model. Each pathway uncovers unique neural circuitry patterns linked to neurological diseases and cognitive functions. The LM pathway offers interpretability through the weights assigned to the input features—vectorized upper triangular portion of the brain network matrix, while the GAT pathway provides interpretability via attention maps averaged across test samples.
To help understand the circuitry patterns captured by two pathways, we map ROIs into six commonly recognized neural systems based on structural and functional roles [44]: the Somatomotor Network (SM), Default Mode Network (DMN), Ventral Salience Network (VS), Central Executive Network (CE), Dorsal Salience Network (DS), and Visual Network (Vis). For this analysis, we focus on the ABCD dataset due to its extensive coverage of cognitive and behavioral measures across a large population, making it particularly suitable for a detailed interpretability study.
@@
Important Connections (Edge-level): To explore the relationship between functional brain networks and cognitive abilities, we begin by visualizing the interpretability results through heatmaps. Fig. 3(3(a)) presents the attention scores from the GAT pathway. It addresses prominent within-system connections, indicated by the diagonal blocks. The cross-system interactions between DMN-CE and Vis-SM also stand out. In contrast, Fig. 3(3(b)) displays the weights from the LM pathway. This heatmap includes both positive (red) and negative (blue) weights, reflecting the direction and magnitude of each connection’s impact on the prediction. Compared to the attention map, the weights are more uniformly distributed, indicating that the LM pathway captures broader connections beyond local interactions.
To complement the global views from the heatmaps, we focus on the top $0.1\%$ of the most significant connections (65 in total). Fig. 3(3(c)) shows the chordplot from the GAT pathway, which emphasizes within-system connectivity, especially within the DMN, SM, and CE systems. There are fewer cross-module connections, most of which involve the DMN. Additionally, there is a noticeable symmetry in the chordplot, with many connections linking the same ROI across the left and right hemispheres, underscoring GAT’s focus on both local and bilateral brain network organizations. In contrast, Fig. 3(3(d)) presents the chordplot from the LM pathway, where connections are more evenly spread across both within-system and cross-system interactions, with a balanced distribution of positive and negative weights.
(a) Heatmap of Attention from GAT Pathway
(b) Heatmap of Weights from LM Pathway
(c) Chordplot of Attention from GAT pathway
(d) Chordplot of Weights from LM pathway
@@
(a) Attention Scores
(b) Positive Weights
(c) Negative Weights
Figure 4: Visualization of the top 20 salient ROIs identified by two pathways in the ABCD dataset. Color intensity represents the importance of each ROI, with darker shades indicating lower importance and brighter shades indicating higher importance. For the GAT pathway, node importance is calculated by summing the attention scores across all its connections, corresponding to each row in the attention heatmap. The identified salient ROIs are primarily concentrated in the Central Executive Network (CE) and Default Mode Network (DMN), with key regions located in the frontal and temporal lobe. For the LM pathway, node importance is calculated separately for positive and negative weights. Salient ROIs identified from the LM pathway’s positive weights are concentrated in the Ventral Salience Network (VS) and Somatomotor Network (SM), predominantly in the temporal lobe, while negative weights highlight regions within the DMN, particularly in the frontal and parietal lobes.
Salient ROIs (Node-level): For the GAT pathway, node importance is determined by summing the attention scores across all connections per node, while for the LM pathway, positive and negative weights are considered separately. We select the top 20 ROIs based on node importance and visualize their locations using the BrainNet Viewer tool [45] in Fig. 4.
@@
(a) Degree Distribution of GAT pathway Subgraph
(b) Degree Distribution of LM pathway Subgraph
(c) Numerical Graph Properties
@@
Graph Properties of the Highlighted Brain Connectivity (Subgraph-level): We now shift our attention to subgraph-level structural characteristics of the brain connectivities identified as important by our model. Specifically, subgraphs are constructed using the top 5% of edges from the GAT and LM pathways respectively, and key graph properties [8]—degree distribution, clustering coefficient, modularity, average shortest path length, global efficiency, small-worldness, and assortativity—are computed to compare the brain connectivities emphasized by each pathway.
The degree distribution, as shown in Fig. 5(5(a)) and 5(5(b)), reveals distinct patterns between the GAT-induced and LM-induced subgraphs. The GAT subgraph displays a more skewed distribution, indicating the presence of a few highly connected hubs, which are characteristics of scale-free networks. While the LM subgraph exhibits a broader distribution with a higher concentration of nodes having intermediate degrees, suggesting a normal distribution of connectivity.
Other graph properties are shown in Fig. 5(5(c)). the GAT subgraph exhibits higher clustering coefficients and modularity, pointing to a graph structure that favors tightly-knit communities and clear modular divisions. The LM subgraph shows shorter average path lengths and higher global efficiency, indicative of an efficient network. Besides, the assortativity coefficient is higher in the LM subgraph, reflecting a stronger tendency for connections between similar nodes. Notably, the small-worldness metric, which reflects a balance between local clustering and global integration, is significantly higher in the GAT subgraph.
Overall, the dual-pathway model demonstrates distinct yet complementary neural circuitry patterns across both pathways. The GAT pathway highlights prominent within-system connectivity and reveals a small-world structure with a strong local clustering phenomenon. While the LM pathway shows more evenly distributed connections and a network structure emphasizing global efficiency and cross-system interactions. Together, these pathways offer a more holistic view of brain organization. The combined insights reinforce the model’s potential to capture diverse neural mechanisms underlying cognitive and behavioral outcomes.

### v2

@@ Abstract
Graph deep learning models, a class of AI-driven approaches employing a message aggregation mechanism, have gained popularity for analyzing the functional brain connectome in neuroimaging. However, their actual effectiveness remains unclear. In this study, we re-examine graph deep learning versus classical machine learning models based on four large-scale neuroimaging studies. Surprisingly, we find that the message aggregation mechanism, a hallmark of graph deep learning models, does not help with predictive performance as typically assumed, but rather consistently degrades it. To address this issue, we propose a hybrid model combining a linear model with a graph attention network through dual pathways, achieving robust predictions and enhanced interpretability by revealing both localized and global neural connectivity patterns. Our findings urge caution in adopting complex deep learning models for functional brain connectome analysis, emphasizing the need for rigorous experimental designs to establish tangible performance gains and perhaps more importantly, to pursue improvements in model interpretability.
@@ 1 Introduction
Brain connectome analysis has captivated researchers for its potential to unravel brain organization and predict clinical outcomes [wang2019hierarchical, long_range1, braingnn, bnt]. Functional magnetic resonance imaging (fMRI) plays a crucial role in this field by capturing brain activity through blood-oxygen-level-dependent (BOLD) signals. These signals facilitate the construction of functional brain networks, where nodes correspond to specific brain regions of interest (ROIs) and edges represent pair-wise statistical relationships, typically correlations, between the BOLD signals of these regions [smith2011network, simpson2013analyzing, edge_centric, complex_measures]. Studying the functional connectome can enhance our understanding of brain functions [dmn_ce2, vim_sm2], enabling us to identify neural circuitry patterns associated with neurological conditions [vim_sm3, pos_neg2] and develop improved diagnostic and therapeutic strategies [sm1, sym2, medial_temp2, mood_disorder].
@@
Machine learning (ML) models have long been employed in functional brain connectome analysis, showing strong robustness and interpretability [pos_neg3, pos_neg2, finn2015functional, shen2017using, cui2018effect, PERVAIZ2020116604, cmep]. These models use the correlation strength between ROIs as predictive features, without considering the underlying graph structure. As AI increasingly permeates healthcare and neuroscience, graph deep learning (GDL) models like graph neural networks (GNNs) [braingnn, braingb, ibgnn, fbnetgen, neurograph] and graph transformers (GTs) [bnt, gated, community_bnt] have recently gained popularity in this field. They explicitly model brain networks as graphs, aggregating messages among nodes (ROIs) along edges (functional connections) [mpnn]. Among various node features, the connection profile—rows of the functional connection matrix—is the most common choice [braingb, fbnetgen, bnt, ibgnn, braingnn, neurograph, gated, community_bnt, hypergraph]. Despite significant efforts to develop GDL models for brain connectome analysis, there is unfortunately no rigorous study and clear understanding of the genuine effectiveness of these advanced models over classical ML models, particularly regarding their unique message aggregation mechanisms.
@@
In this work, we conduct a comprehensive evaluation of GDL against classical ML models, across four fMRI-based brain connectome datasets: ABIDE [abide], PNC [pnc], HCP [hcp], and ABCD [abcd]. These datasets cover a diverse population and clinical prediction tasks, from demographic attributes to cognitive abilities and neurological disorders. The experimental results are surprising: (1) classical ML models, such as logistic regression and kernel ridge regression, along with simple multi-layer perceptrons (MLPs), often match or even exceed the predictive performance of more complex GDL models; and (2) the message aggregation mechanisms unique to GDL models and believed to be a key technical advantage, appear to have a consistent detrimental effect on prediction performance within the context of functional brain networks. These results challenge prevailing assumptions about the superiority of GDL methods in modeling brain connectome.
@@
Motivated by the effectiveness of classical models and the limitations of GDL, we introduce a simple yet effective model featuring a novel dual-pathway design, which synergizes linear modeling (LM) with a graph attention network (GAT) [gat]. The LM pathway efficiently captures global connectivity patterns from whole connectivity matrices, while the GAT pathway adaptively extracts local graph structures using embeddings of BOLD time series as node features. Despite its simple architecture, experimental results demonstrate that our model consistently achieves robust and competitive performance across diverse datasets, rivaling or surpassing the best-performing baselines. Beyond predictive performance, the model offers more holistic and complementary interpretability through its dual-pathway design: the GAT pathway is particularly effective at highlighting localized subnetworks and functional hubs, offering insights into modularized brain organization, while the LM pathway emphasizes global connectivity patterns and network-wide efficiency. The patterns identified not only align well with established neuroscientific knowledge but also reveal specific functional connections, salient ROIs, and subnetworks associated with cognitive processes, potentially providing finer-grained insights into neural mechanisms.
@@
The findings of this study suggest caution with aggregation-based GDL models, particularly when connection profiles are used as node features, where aggregation shows clear negative effects. They underscore the importance of thoroughly benchmarking new models in future brain connectome research. Meanwhile, we advocate for future research to prioritize interpretability and contextual knowledge for scientific or clinical applications over simply pursuing prediction accuracy.
@@
Figure 1: Performance comparison across different baselines on four datasets. For the ABIDE and PNC datasets, the prediction tasks are binary classification for autism disease and biological gender, respectively, evaluated using AUROC scores. For the HCP and ABCD datasets, the prediction task is fluid intelligence score prediction, a regression task evaluated using the Pearson correlation between predicted and actual values. The baselines are grouped into three categories, distinguished by color and hatching patterns: green box for graph deep learning, blue box for non‑graph deep learning, and red box for classical machine learning models. Models highlighted with bold borders, colored hatching and yellow highlight band represent those with the best predictive performance among all the baselines. There is no significant difference ($p>0.05$) in performance among these highlighted models, showing that simpler models such as Logistic Regression, MLP, ElasticNet, and Kernel Ridge Regression perform comparably to the most advanced graph deep learning models. The significance bars marked with asterisks (*) indicate that the performance of non-graph deep learning models is significantly better than that of the graph deep learning models displayed to the left ($p<0.05$). The results show that simple models can match or even exceed the predictive performance of more complex GDL models.
@@ 2 Results
To address the ambiguity regarding the effectiveness of complex GDL models, we conduct extensive experiments across four fMRI-based brain network datasets with various clinical prediction tasks, ranging from demographics to cognitive ability and neural disorders: ABIDE [abide] for autism disease classification, PNC [pnc] for gender prediction, and both HCP [hcp] and ABCD [abcd] for fluid intelligence prediction. Our benchmark spans three types of models: GDL, non-graph deep learning, and classical ML models.
Fig. 1 illustrates the prediction performance of each baseline across various datasets. Simpler models often match or outperform more advanced GDL models. Specifically, for classification tasks on ABIDE and PNC datasets, classical models (e.g., Logistic Regression) and MLP-based models consistently outperform most GDL models, with performance comparable to state-of-the-art models like NeuroGraph and BrainNetTF. Similarly, for regression tasks on HCP and ABCD datasets, classical ML models such as ElasticNet [PERVAIZ2020116604] and Kernel Ridge Regression [pos_neg2] demonstrate superior performance over most GDL models, and perform comparably to BrainNetTF [bnt], one of the most advanced aggregation-based GDL models. These findings challenge the prevailing belief that graph-based modeling of brain connectome is superior regarding the prediction of clinical outcomes.
@@
In GDL models, the message aggregation is crucial for combining node features from neighboring nodes to form new node representations. However, as shown in Fig. 1, most aggregation-based GDL models tend to underperform classical ML models like Logistic Regression, SVM/SVR [hastie2009elements] and Kernel Ridge Regression, and simpler MLP-based models. This raises the question about the genuine effectiveness of the aggregation operations in brain connectome analysis.
@@
To further investigate the impact of aggregation, we vary the graph density by retaining different percentages of the top $K\%$ edges in the graphs, so as to explicitly control the extent of aggregation. Notably, with $K=0$, the graphs have no edges between ROIs, so node features are transformed independently, eliminating any aggregation effects. To align with existing studies, the original settings of each GDL model are followed for extracting the top $K\%$ edges. Specifically, for models like GCN [gcn], GAT [gat], GIN [gin], GraphSage [graphsgae], NeuroGraph [neurograph], and BrainGNN [braingnn], only positive correlations between ROIs are considered as potential edges. This approach adheres to the common assumption that positive correlations are typically more informative in the context of functional brain networks [braingnn, neurograph]. Conversely, for BrainGB [braingb] and BrainNetTF [bnt], both positive and negative correlations are incorporated, following the methodologies described in the original papers.
@@
The left panels of the Fig. 2 show that models like GCN, GAT, GIN, and GraphSage consistently exhibit a decline in prediction performance as the graph density increases, indicating that message aggregation does not enhance prediction and more aggregation can lead to more performance degradation. Similarly, BrainGB and BrainGNN demonstrate optimal performance with lower density, aligning with the notion that message aggregation may not be helpful in brain connectome analysis tasks.
@@
Unlike other GDL models, BrainNetTF and NeuroGraph incorporate not only message aggregation but also residual connections, which integrate the untransformed input brain connectivity with the transformed hidden embeddings. The right panels of the Fig. 2 indicate that the performance of NeuroGraph and BrainNetTF remains relatively stable across varying graph density. However, when residuals are removed, as in the variants NeuroGraph w/o Res and BrainNetTF w/o Res, performance declines as density increases, consistent with other GDL models. This further underscores the point that message aggregation alone impairs predictive performance in brain connectome analysis.
Figure 3: Relationship between node representation smoothness and predictive performance. Each panel shows the relationship between model performance (y-axis) and the mean cosine similarity of the last-layer node embeddings (x-axis), which quantifies the smoothness of node representations. Results are presented for four datasets and three representative GDL models (GCN, GAT, and BrainNetTF without residual connections). Each point corresponds to the average over 10 independent runs under a specific graph density, with the adjacent number indicating the corresponding density level. Models are distinguished by color and marker type. The Pearson correlation coefficient between predictive performance and cosine similarity and its $p$-value are reported in the legend. Across all datasets and models, node smoothness and predictive performance exhibit a strong and statistically significant negative correlation (Pearson $r<0$, all $p<0.05$). As graph density increases, message aggregation leads to progressively smoother (more homogeneous) node representations, which coincide with decreased predictive accuracy. This indicates a manifestation of the over-smoothing phenomenon that underlies the limited effectiveness of GDLs in functional brain connectome analysis.
To further probe the mechanism behind the observed degradation in predictive performance, we conduct an additional analysis focusing on the smoothness of the learned node features by GDL models. Specifically, we examine three representative GDL models—GCN, GAT, and BrainNetTF (w/o residuals), which together capture the main architectural paradigms of graph convolutional networks and graph transformers. The residual connections in BrainNetTF are removed in this analysis to better isolate the effect of message aggregation itself. For each model and dataset, we compute the mean cosine similarity of the last-layer node embeddings (i.e., the representations directly used for prediction) across varying graph densities. This metric is widely adopted to measure the smoothness of representation [chen2020measuring], where higher values indicate that node embeddings become more homogeneous. As shown in Fig. 3, across all four datasets and three representative GDL models, we consistently observe that as graph density increases, message aggregation generally leads to progressively smoother node features, accompanied by a decline in predictive accuracy. Moreover, the analysis reveals a strong and statistically significant negative correlation between node smoothness and model performance (Pearson $r<0$ and $p<0.05$). This consistent inverse relationship provides quantitative evidence that, in the context of functional brain connectome analysis, message aggregation tends to make node features converge and lose discriminability, a manifestation of the over-smoothing phenomenon that underlies the limited effectiveness of GDL models.
Table 1: Effective rank of input node-feature matrices. For each dataset, the table reports the mean effective rank of the input node-feature matrices (averaged across subjects) together with the total feature dimensionality. The Rank/Dim ratio quantifies the proportion of independent variation within the input features. Across all datasets, the effective ranks are notably low (7–38% of the feature dimensionality), indicating that the connection-profile-based node features lie in a low-dimensional subspace and contain substantial redundancy and correlation among nodes. This low-rank property contributes to the tendency of graph message aggregation to induce over-smoothing in learned representations.
Dataset Node Feature Dim Effective Rank Rank/Dim Ratio
ABIDE 200 26.02 0.13
PNC 264 19.35 0.07
HCP 132 49.94 0.38
ABCD 360 33.84 0.09
In functional brain connectome analysis, GDL models typically use the functional connection profile as the node feature, which has become a standard practice. Each node feature corresponds to a row of the functional connectivity matrix and thus inherently encodes the to all other nodes in the network. Consequently, every node feature already carries global information, leading to substantial correlation and redundancy among nodes even before any message aggregation. To quantify this redundancy, we examine the effective rank [roth2024rank] of the input node feature matrices for each dataset. As shown in Table 1, the average effective ranks are notably low relative to the input dimensionality (7–38% of feature dimensions). Given such low-rank and highly correlated inputs, message aggregation—which essentially acts as a low-pass filter over the graph [li2018deeper], further amplifies shared components among nodes and dilutes distinctive variations, making the learned representations more homogeneous and thus more prone to over-smoothing.
Collectively, these findings indicate that in functional connectome analysis, the inherent low-rank property of brain connectivity features makes GDLs vulnerable to over-smoothing under message aggregation, which in turn degrades their predictive capability. This also explains why architectures with residual or skip connections (e.g., NeuroGraph [neurograph] and BrainNetTF [bnt]) remain more stable performance, as they preserve variance from the original, unaggregated node features and thereby mitigating over-smoothing.
@@
Previous experiments have shown that simple linear models like Logistic Regression and ElasticNet, often outperform other baselines. However, relying solely on linear models overlooks graph structure information inherent in brain networks, which is essential for understanding brain function and cognitive processes [complex_measures, clustering_coef1]. To incorporate this information, it is natural to consider using graph models. However, Fig. 2 reveals that using connection profiles as node features diminishes the effectiveness of aggregation operations within graph models.
To address these challenges, we propose a dual-pathway model that integrates the strengths of both linear non-graph and aggregation-based graph models. The model consists of two complementary pathways. The first pathway retains the simplicity and effectiveness of linear models, which uses the flattened upper triangular part of the functional brain network matrix as input. The second pathway utilizes graph structure by embedding BOLD time series, rather than connection profiles, as node features. These BOLD embeddings, which are then processed by a graph attention network (GAT), are obtained through a 1D-CNN encoder that captures local temporal dynamics of each ROI, providing more diverse and less redundant inputs for graph aggregation. Table 2 shows that this dual-pathway model outperforms baselines on the ABCD and PNC datasets, matches performance on ABIDE, and exhibits slight underperformance on HCP. Overall, this dual-pathway model achieves robust and competitive performance across four datasets.
@@
The proposed dual-pathway model integrates a non-aggregation linear model (LM) and an aggregation-based graph attention network (GAT). Each pathway uncovers unique neural circuitry patterns linked to brain functions. The LM pathway offers interpretability through the weights assigned to the input features—vectorized upper triangular portion of the brain network matrix, while the GAT pathway provides interpretability via attention maps averaged across test samples.
In this section, we use the ABCD dataset as a representative case study to demonstrate the interpretability of the dual-pathway model. ABCD is selected for its large sample size (over 7,000 participants), which supports statistical stability and robust estimation of model-derived circuitry patterns [marek2022reproducible]. It also employs a high-resolution 360-region parcellation, providing broad spatial coverage of ROIs and sufficient granularity to identify interpretable neural systems. To help understand the circuitry patterns captured by two pathways, we map ROIs into six commonly recognized neural systems based on structural and functional roles [functional_module]: the Somatomotor Network (SM), Default Mode Network (DMN), Ventral Salience Network (VS), Central Executive Network (CE), Dorsal Salience Network (DS), and Visual Network (Vis).
@@
Important Connections (Edge-level): To explore the relationship between functional brain networks and cognitive abilities, we begin by visualizing the interpretability results through heatmaps. Fig. 4(a) presents the attention scores from the GAT pathway. It addresses prominent within-system connections, indicated by the diagonal blocks. The cross-system interactions between DMN-CE and Vis-SM also stand out. In contrast, Fig. 4(b) displays the weights from the LM pathway. This heatmap includes both positive (red) and negative (blue) weights, reflecting the direction and magnitude of each connection’s impact on the prediction. Compared to the attention map, the weights are more uniformly distributed, indicating that the LM pathway captures broader connections beyond local interactions.
To complement the global views from the heatmaps, we focus on the top $0.1\%$ of the most significant connections (65 in total). Fig. 4(c) shows the chordplot from the GAT pathway, which emphasizes within-system connectivity, especially within the DMN, SM, and CE systems. There are fewer cross-module connections, most of which involve the DMN. Additionally, there is a noticeable symmetry in the chordplot, with many connections linking the same ROI across the left and right hemispheres, underscoring GAT’s focus on both local and bilateral brain network organizations. In contrast, Fig. 4(d) presents the chordplot from the LM pathway, where connections are more evenly spread across both within-system and cross-system interactions, with a balanced distribution of positive and negative weights.
@@
Figure 5: Visualization of the top 20 salient ROIs identified by two pathways in the ABCD dataset. Color intensity represents the importance of each ROI, with darker shades indicating lower importance and brighter shades indicating higher importance. (a) For the GAT pathway, node importance is calculated by summing the attention scores across all its connections, corresponding to each row in the attention heatmap. The identified salient ROIs are primarily concentrated in the Central Executive Network (CE) and Default Mode Network (DMN), with key regions located in the frontal and temporal lobe. (b)-(c) For the LM pathway, node importance is calculated separately for positive and negative weights. Salient ROIs identified from the LM pathway’s positive weights are concentrated in the Ventral Salience Network (VS) and Somatomotor Network (SM), predominantly in the temporal lobe, while negative weights highlight regions within the DMN, particularly in the frontal and parietal lobes.
Salient ROIs (Node-level): For the GAT pathway, node importance is determined by summing the attention scores across all connections per node, while for the LM pathway, positive and negative weights are considered separately. We select the top 20 ROIs based on node importance and visualize their locations using the BrainNet Viewer tool [brainnetviewer] in Fig. 5.
@@
Graph Properties of the Highlighted Brain Connectivity (Subgraph-level): We now shift our attention to subgraph-level structural characteristics of the brain connectivities identified as important by our model. Specifically, subgraphs are constructed using the top 5% of edges identified as important from the GAT and LM pathways respectively, and key graph properties [complex_measures]—degree distribution, clustering coefficient, modularity, average shortest path length, global efficiency, small-worldness, and assortativity—are computed to compare the brain connectivities emphasized by each pathway.
The degree distribution, as shown in Fig. 6(a) and 6(b), reveals distinct patterns between the GAT-induced and LM-induced subgraphs. The GAT subgraph displays a more skewed distribution, indicating the presence of a few highly connected hubs, which are characteristics of scale-free networks. While the LM subgraph exhibits a broader distribution with a higher concentration of nodes having intermediate degrees, suggesting a normal distribution of connectivity.
Other graph properties are shown in Fig. 6(c). the GAT subgraph exhibits higher clustering coefficients and modularity, pointing to a graph structure that favors tightly-knit communities and clear modular divisions. The LM subgraph shows shorter average path lengths and higher global efficiency, indicative of an efficient network. Besides, the assortativity coefficient is higher in the LM subgraph, reflecting a stronger tendency for connections between similar nodes. Notably, the small-worldness metric, which reflects a balance between local clustering and global integration, is significantly higher in the GAT subgraph.
The dual-pathway model demonstrates distinct yet complementary neural circuitry patterns across both pathways. The GAT pathway highlights prominent within-system connectivity and reveals a small-world structure with a strong local clustering phenomenon. The LM pathway shows more evenly distributed connections and a network structure emphasizing global efficiency and cross-system interactions. Together, these pathways offer a more holistic view of brain organization. The combined insights reinforce the model’s potential to capture diverse neural mechanisms underlying cognitive and behavioral outcomes.